OpenStudy (anonymous):

Point R is located at (-5, 1), and S is located at (4, -6). Find the coordinates of the point that lies halfway between R and S. (5 points) Select one: a. (-2, -1) b. (-3, 5) c. (-0.5, -2.5) d. (-4.5, 3.5)

hartnn (hartnn):

"point that lies halfway" is a mid-point

hartnn (hartnn):

so, use the mid-point formula The co-ordinates of midpoint(x,y) on the line segment with end-points (x1,y1) and (x2,y2) is just the average of co-ordinates, given by : \(\large x=\frac{x_1+x_2}{2}\) \(\large y=\frac{y_1+y_2}{2}\) now,just put the values and find the co-ordinates of midpoint (x,y).

OpenStudy (anonymous):

ok i dont get this

hartnn (hartnn):

just take the average of the corresponding co-ordinates! (-5+4)/2 = ...? (1-6)/2 = ...?